OpenStudy (anonymous):

Solve and graph the absolute value inequality: |2x + 1| ≤ 5. number line with closed dots on -3 and 2 with shading going in the opposite directions. number line with closed dots on -3 and 2 with shading in between. number line with open dots on -3 and 2 with shading in between. number line with closed dots on -2 and 2 with shading in between.

jagr2713 (jagr2713):

Well its ≤ so its a closed dot so we scratch out C

OpenStudy (anonymous):

Well that helps a bit

jagr2713 (jagr2713):

First the x is not negative so it will still go the same way. So its pointing left, so it didnt change so we can scratch A and when you subtract 5-1=4 and 4 divided by 2=2 and its still going to the left and 5+1=6 6 and 6 divided by 2=3 and going to left so it isnt opposite so it would be B in my opinion

OpenStudy (cwrw238):

2x + 1 <= 5 2x <= 4 x <= 2 also 2x + 1 >= -5 2x > -6 x >= -3

OpenStudy (cwrw238):

B is correct

OpenStudy (cwrw238):